OpenStudy (isbella):

1)Sketch the vector field F⃗(x,y)=5i−4j and calculate the line integral of F⃗ along the line segment from (−5,−1) to (−3,0). help anyone i dont know how to do this

OpenStudy (phi):

the vector field is just an "arrow" in the direction 5 over, 4 down at every point. the line integral is \[ \int_C \vec{F}\cdot d\vec{r} \] where \(\vec{r}\) is the position vector that describes the curve. in this case r⃗ is a line given by the equation y = ½ x + 3/2 in other words, \[ (\vec{r}= <x, ½ x + 3/2> \] find the derivative of r⃗ with respect to x, and calculate \[ \int \vec{F}\cdot \frac{d\vec{r}}{dx} \ dx \]

OpenStudy (phi):

you should get \[ \vec{F}= <5,-4> \\\frac{d\vec{r}}{dx} = <1,\frac{1}{2}> \] and the integral \[ \int_{-5}^{-3} <5,-4> \cdot < 1, \frac{1}{2}>\ dx \\ \int_{-5}^{-3} 3 \ dx = 3 x \bigg|_{-5}^{-3} = -9 +15 = 6 \]